roduct Costs using Activity Rates

Atlas Enterprises Inc. manufactures elliptical exercise machines and treadmills. The products are produced in its Fabrication and Assembly production departments. In addition to production activities, several other activities are required to produce the two products. These activities and their associated activity rates are as follows:

Activity Activity Rate
Fabrication $25 per machine hour
Assembly $15 per direct labor hour
Setup $53 per setup
Inspecting $26 per inspection
Production scheduling $16 per production order
Purchasing $13 per purchase order

The activity-base usage quantities and units produced for each product were as follows:

Activity Base Elliptical Machines Treadmill
Machine hours 1,755 1,035
Direct labor hours 443 173
Setups 46 14
Inspections 740 444
Production orders 64 13
Purchase orders 173 106
Units produced 281 188

Use the activity rate and usage information to calculate the total activity cost and activity cost per unit for each product. If required, round the per unit answers to the nearest cent.

Total Activity Cost Activity Cost Per Unit
Elliptical Machines $ $
Treadmill $ $

Contribution Margin, Break-Even Sales, Cost-Volume-Profit Chart, Margin of Safety, and Operating Leverage

Belmain Co. expects to maintain the same inventories at the end of 20Y7 as at the beginning of the year. The total of all production costs for the year is therefore assumed to be equal to the cost of goods sold. With this in mind, the various department heads were asked to submit estimates of the costs for their departments during the year. A summary report of these estimates is as follows:

Estimated
Fixed Cost
Estimated Variable Cost
(per unit sold)
Production costs:
Direct materials $24
Direct labor 16
Factory overhead $128,400 12
Selling expenses:
Sales salaries and commissions 26,700 5
Advertising 9,000
Travel 2,000
Miscellaneous selling expense 2,200 4
Administrative expenses:
Office and officers' salaries 26,100
Supplies 3,200 2
Miscellaneous administrative expense 3,040 3
Total $200,640 $66

It is expected that 6,840 units will be sold at a price of $132 a unit. Maximum sales within the relevant range are 9,000 units.

Required:

1. Prepare an estimated income statement for 20Y7.

Belmain Co.
Estimated Income Statement
For the Year Ended December 31, 20Y7
$
Cost of goods sold:
$
Total cost of goods sold
Gross profit $
Expenses:
Selling expenses:
$
Total selling expenses $
Administrative expenses:
$
Total administrative expenses
Total expenses
Income from operations $

2. What is the expected contribution margin ratio? Round to the nearest whole percent.
%

3. Determine the break-even sales in units and dollars.

Units units
Dollars units

4. Construct a cost-volume-profit chart on your own paper. What is the break-even sales?
$

5. What is the expected margin of safety in dollars and as a percentage of sales?

Dollars: $
Percentage: (Round to the nearest whole percent.) %

6. Determine the operating leverage. Round to one decimal place.
Answer #1

Calculate total activity cost and cost per unit

Elliptical Machines Treadmill
Fabrication 1755*25 = 43875 1035*25 = 25875
Assembly 443*15 = 6645 173*15 = 2595
Setup 46*53 = 2438 14*53 = 742
Inspecting 740*26 = 19240 444*26 = 11544
Production scheduling 64*16 = 1024 13*16 = 208
Purchasing 173*13 = 2249 106*13 = 1378
Total activity Cost 75471 42342
Units 281 188
Unit activity cost 268.58 225.22